The area of a square is 12 more than its perimeter. Find the length of a side.
Answer by mananth(16946)


let side be x
Area of square =Side * Side = x^2
Perimeter of square = 4*side= 4x
Area is 12 times the perimeter
x^2 = 12 *4*x
x^2/x= 48
x=48
the side is 48